OpenStudy (anonymous):

yes or no The lengths of the sides of a triangle are 4, 5, 6. Can the triangle be a right triangle?

OpenStudy (anonymous):

4^2 + 5^2 = 16 + 25 = 41 which is NOT equal to 6^2 Since It cannot fulfill Pythagoras theorem, it cannot be a right triangle. NO.